Please help me with this word problem: a glider flies 8 miles south from the airport and then 15 miles east. then it flies in a straight line back to the airport. What was the distance of the gliders last leg back to the airport?
Found 2 solutions by unlockmath, ammuann123:
Answer by unlockmath(1688)

Hello,
We're dealing with a right triangle so we can apply the following formula:
A^2 + B^2 = C^2 (C will be the leg back) Now we can plug in the numbers.
8^2 + 15^2 = C^2 Do the calculations will result in:
64 + 225 = c^2 Combine like terms:
289 = c^2 Square root each side will give us:
c = 17
There we go. the answer is:
17 Miles
Let me know if this makes sense to you.
Check out my book at www.math-unlock.com
RJ Toftness

Answer by ammuann123(2)

You must use pythagorean thoerem.
x^2= (15)^2 + (8)^2
x^2= 289
x= Sqrt(289)
x= 17 miles